OpenStudy (anonymous):

sarah is twice as old as john. six years ao, sarah was 4 times as old as john was then. how old is john now?

terenzreignz (terenzreignz):

First step in algebraic word problems is to represent values :) If we let x be the age of John now, what is the age of Sarah?

OpenStudy (anonymous):

2x?

terenzreignz (terenzreignz):

Good :) Okay, so what would be John's age SIX YEARS AGO ?

OpenStudy (anonymous):

2x-6? :)

terenzreignz (terenzreignz):

uh-uh, that's ROSE's age six years ago. What about John?

terenzreignz (terenzreignz):

Remember, John's age now is x.

terenzreignz (terenzreignz):

omg fail

OpenStudy (anonymous):

um well idk

terenzreignz (terenzreignz):

forget I said anything... that's Sarah

terenzreignz (terenzreignz):

Sarah's age 6 years ago was 2x - 6

OpenStudy (anonymous):

lol XD i was so confused

terenzreignz (terenzreignz):

Sorry, rosie, I was putting you in the problem XD

OpenStudy (texaschic101):

lol

terenzreignz (terenzreignz):

gah

terenzreignz (terenzreignz):

But anyway, John's age now is x, Sarah's age now is 2x. Six years ago, their ages should be...?

OpenStudy (anonymous):

so sarahs age is six yrs ago was 2x-6

terenzreignz (terenzreignz):

yup. What about John?

OpenStudy (anonymous):

and john is 4x-6?

terenzreignz (terenzreignz):

No... John's age now is x. so six years ago, it should be...? (if you're 20 now, six years ago it would be 14...)

OpenStudy (anonymous):

x-6?

terenzreignz (terenzreignz):

Yes, precisely :D

terenzreignz (terenzreignz):

John's age six years ago = x - 6 Sarah's age six years ago = 2x - 6 Now, to quote your problem: Six years ago, Sarah was 4 times as old as John was then. How do you translate that in math terms?

OpenStudy (anonymous):

4(2x-6)(x-6)?

OpenStudy (anonymous):

or is it 4(x-6)

terenzreignz (terenzreignz):

Okay, I'll translate, you solve... :) Trust me, the only difficult thing in these word problems is finding what problem to solve... when you get it, it's usually easy :) Now... Six years ago... (this only means we'll be using their ages six years ago) Sarah 2x - 6 was 2x - 6 = 4 times 2x - 6 = 4() as old as John was then 2x - 6 = 4(x - 6) Now just solve for x... which we know to be John's age now :D

OpenStudy (anonymous):

oh my god thank you so much. this is from the book for the sats. i was studying so yea :)

terenzreignz (terenzreignz):

Thank me by telling me your answer :3

OpenStudy (anonymous):

ok so i got 9. is that correct?

terenzreignz (terenzreignz):

Bingo :)

OpenStudy (anonymous):

yay finally got it. thanks so much :)

terenzreignz (terenzreignz):

No problem :)